### v1

# Static or Dynamic: Towards Query-Adaptive Token Selection for Video Question Answering

Abstract
Video question answering benefits from the rich information available in videos, enabling a wide range of applications. However, the large volume of tokens generated from longer videos presents significant challenges to memory efficiency and model performance. To alleviate this issue, existing works propose to compress video inputs, but usually overlooking the varying importance of static and dynamic information across different queries, leading to inefficient token usage within limited budgets. To tackle this, we propose a novel token selection strategy, explore-then-select, that adaptively adjust static and dynamic information needed based on question requirements. Our framework first explores different token allocations between static frames, which preserve spatial details, and dynamic frames, which capture temporal changes. Next, it employs a query-aware attention-based metric to select the optimal token combination without model updates. Our proposed framework is plug-and-play that can be seamlessly integrated within diverse video-language models. Extensive experiments show that our method achieves significant performance improvements (up to 5.8%) among various video question answering benchmarks.
Static or Dynamic: Towards Query-Adaptive Token Selection for Video Question Answering
Yumeng Shi  Quanyu Long  Wenya Wang Nanyang Technological University yumeng001@e.ntu.edu.sgquanyu001@e.ntu.edu.sgwangwy@ntu.edu.sg
1 Introduction
Video Question Answering (VideoQA) has broad applications across various fields (Mogrovejo and Solorio, 2024; Zhang et al., 2024a). Compared to text, videos provide more intuitive and dynamic information, delivering richer context and details by combining visual and temporal elements. Current research primarily leverages powerful large language models to build video-language models (VideoLMs) (Lin et al., 2023; Zhang et al., 2024b), significantly enhancing AI performance in VideoQA tasks. However, the extensive visual information in long videos leads to a dramatic increase in token counts. For instance, if one frame generates 196 tokens (Li et al., 2024a), a 5-minute video sampled at 1fps would produce nearly 60,000 tokens, posing significant challenges to memory requirements and model capabilities.
Figure 1: Different question types vary in their dependence on static and dynamic information in videos. For example, Question 2 relies on fine-grained dynamic information, while Question 1 and 3 only require static frames. The frames needed to answer the questions are highlighted with corresponding colored boxes.
Given the strict token limitations in practical VideoLM deployments, effectively representing essential video information requires a careful allocation between static and dynamic content. Static information, which refers to the visual content within individual frames, is crucial for questions like object recognition, where spatial details dominate. In contrast, dynamic information captures temporal changes and motion patterns across consecutive frames, which are essential for understanding actions or events. Figure 1 illustrates different types of questions, which vary in their reliance on static and dynamic information. Considering these varying dependencies, the challenge lies in optimizing the allocation of limited tokens to preserve the most relevant aspects of both static and dynamic information, depending on specific question requirements. Although existing studies Shen et al. (2024); Nie et al. (2024) have explored token compression through changing frame sampling rates or intra-frame downsampling, they fail to address the varying dependencies on static and dynamic information across different question types.
To achieve an effective allocation between static and dynamic information in video token compression, we propose a novel token selection strategy, Explore-then-Select, that adaptively aligns video tokens with textual queries under a limited token budget. Unlike previous approaches that rely on fixed rules, our strategy autonomously and adaptively combines static and dynamic content based on the nature of the questions (e.g., action description, event sequence, or object recognition), ensuring more precise responses to diverse queries.
Specifically, we categorize video frames into static and dynamic frames. Static frames are fully retained to preserve essential spatial details, such as objects, while dynamic frames are selectively processed, retaining only a subset of tokens to capture critical temporal changes and motion patterns. To optimize token allocation between these two types of frames, Explore-then-Select uses a two-stage process. In the exploration stage, we construct a search space consisting of various combinations of static and dynamic frames, where each combination generates a token subsequence of the same length (due to a fixed token budget). By adjusting the proportion of static and dynamic frames, we can prioritize either static details or dynamic changes based on question requirements. In the selection stage, we evaluate each combination using an query-aware metric derived from the shallow attention layers of VideoLMs. This metric quantifies the alignment between the query and video tokens, enabling us to select the optimal combination to answer the question.
Notably, our framework is training-free, as neither the exploration nor selection processes require model updates. Leveraging its seamless integration with diverse VideoLMs, we demonstrate the effectiveness of our approach on two widely recognized VideoLMs across multiple benchmarks for both long and short videos. Using our framework, models can achieve improvements of up to 5.8%. Our key contributions are summarized as follows:
  • Building on the observation that questions rely differently on static and dynamic video information, we propose a novel Explore-then-Select framework to adaptively and effectively select video tokens reflecting the optimal balance of static and dynamic information under limited token budgets.
  • To address static and dynamic information needs, we design an effective search space of static-dynamic frame combinations. During the selection phase, we employ a query-aware approach, leveraging an attention-based metric to adaptively evaluate candidates and select the optimal combination for each question.
  • We conduct experiments on both long-video and short-video benchmarks, demonstrating the effectiveness of our method. Furthermore, we provide an in-depth analysis of the framework’s characteristics and behavior.
2 Related Work
2.1 Video Language Models
Significant progress has been made in video language model research based on LLMs. These models can be primarily classified into two types: general-purpose vision-language models Team et al. (2024); Chen et al. (2024b); OpenAI (2024); Yao et al. (2024); Ye et al. (2023) and specialized video language models Lin et al. (2023); Zhang et al. (2024c); Li et al. (2024c); Zhang et al. (2025); Liu et al. (2024). In the realm of general-purpose vision-language models, LLaVA-OneVision Li et al. (2024a) unifies single-image, multi-image, and video tasks, while Qwen2-VL Wang et al. (2024) introduces dynamic resolution support and three-dimensional positional encoding for enhanced visual feature capture. In the field of specialized models, VideoChat Li et al. (2023b) focuses on achieving in-depth understanding and interaction with video content. LongVA Zhang et al. (2024b) has successfully extended the context length of language models, transferring their advantages in long-text processing to the video domain.
Method Pre- Training- Video-
Input Free Specific
FastV (Chen et al., 2024a)
ZipVL (He et al., 2024b)
FrameFusion Fu et al. (2024b)
TokenPacker Li et al. (2024b)
VideoStreaming Qian et al. (2025)
SlowFocus Nie et al. (2024)
LongVU Shen et al. (2024)
Ours
Table 1: Feature comparison with existing methods. “Pre-Input” refers to methods which reduce tokens before feeding them into large language models, while “Video-Specific” denotes methods which leverage the unique characteristics of video data.
2.2 Visual Token Compression
Some studies Bolya et al. (2022) focus on compressing visual tokens in video encoders. Among them, the RLT Choudhury et al. (2024) effectively reducing the number of tokens by replacing repeated patches in videos with a single patch. Some works Li et al. (2024b); Qian et al. (2025); Shen et al. (2024); Lan et al. (2024) introduce additional modules for visual token compression. For instance, BLIP-2 Li et al. (2023a) innovatively introduces the Q-Former module, which utilizes learnable query vectors to transform visual features into semantic representations, achieving efficient visual token compression. Furthermore, inspired by KV cache compression Zhang et al. (2023) in long-text processing, some research adopts it in visual token compression He et al. (2024b); Chen et al. (2024a); Fu et al. (2024b). These methods optimize token usage efficiency by setting thresholds based on specific metrics to prune visual tokens.
We compare existing methods in Table 1, noting that training-free approaches mainly compress tokens within the KV cache, reducing FLOPs but not addressing the issue of excessive token input to the large language model. In contrast, methods that reduce tokens in advance typically require training. This paper introduces a novel pre-input, training-free framework for more effective compression, considering query-aware static and dynamic information balancing.
3 Preliminary
In this section, we outline the common inference pipeline of VideoLMs as the setup for our approach. It consists of three key steps, including video frame sampling, visual encoding and embedding, and multimodal inference.
  • Video Frame sampling. Given an input video, $N$ frames are uniformly sampled to form a representation $\bm{V}\in\mathbb{R}^{N\times C\times H_{v}\times W_{v}}$, where $C=3$ denotes the RGB channels, and $H_{v}$ and $W_{v}$ represent the height and width of each frame, respectively.
  • Visual Encoding and Embedding. The sampled frames are decomposed into non-overlapping spatiotemporal patches, which are processed by a vision encoder to extract spatiotemporal features. These features are projected into the language model’s token space via a linear projector, resulting in visual token embeddings $\bm{F}\in\mathbb{R}^{T\times H\times W\times D}$, where $T$ represents the temporal resolution, $H$ and $W$ denote the spatial resolutions, and $D$ is the token embedding dimension.
  • Multimodal Processing. The visual token embeddings $\bm{F}$ are then flattened into a sequence $\bm{T}_{v}\in\mathbb{R}^{THW\times D}$, where the sequence length is $L=T\times H\times W$. The sequence $\bm{T}_{v}$, instruction embeddings $\bm{T}_{i}$, and query embeddings $\bm{T}_{q}$ are concatenated into a unified input $\bm{T}=[\bm{T}_{i},\bm{T}_{v},\bm{T}_{q}]$, where []delimited-[][$\cdot$][ ⋅ ] denotes token concatenation. Finally, VideoLM processes the unified input sequence $\bm{T}$ to generate a textual response to the question.
Figure 2: Overview of our Explore-then-Select framework for token selection. In the exploration stage, we generate different subsequences of different combinations of static and dynamic tokens. In the selection stage, we evaluate their query-aware metrics based on shallow attention layers and select the optimal one for input to LLMs.
4 Method
4.1 Problem Definition
Due to GPU memory and model capability constraints, the number of visual tokens processed during inference is capped at $L_{b}$. The fixed token budget limits frame sampling to a reduced number of frames, resulting in significant loss of rich visual information, particularly in long videos.
In this work, we aim to sample more frames to expand the amount of information we can capture, which generates an excessive number of tokens, leading to a sequence length $L\gg L_{b}$. Then we compress the tokens to meet the token budget, enabling more effective utilization of rich visual information within the limited length.
To meet our goal, we propose a token-efficient framework that automatically and adaptively selects a limited yet informative set of visual tokens by leveraging the textual query’s relevance to both static and dynamic visual information. Our method emphasizes balancing these two types of information, ensuring that the selected tokens maximize their alignment with the query while maintaining memory efficiency.
4.2 Framework Overview
We adopt an Explore-then-Select framework, as illustrated in Figure 2. In the token exploration stage (Section 4.3), we construct a search space of $n$ visual token subsequences, each of length $L_{b}$, where every visual token subsequence reflects a distinct balance of static and dynamic information. In the token selection stage (Section 4.4), we identify the optimal sequence that best aligns with the query requirements. Details will be discussed below.
4.3 Exploration: Search Space Design
This section describes the generation of $n$ token subsequences, each of length $L_{b}$, from a token sequence of length $L$. To balance static and dynamic information in videos based on query requirements, we classify frames into static and dynamic frames. It is worth noting that since some models employ temporal reduction, here we refer to “frames” as $\bm{F}$ and the number of frames is $T$. Based on whether the tokens in a subsequence originate from static or dynamic frames, we divide them into two subsets: static frame tokens $\mathcal{T}_{\text{Static}}$ and dynamic delta tokens $\mathcal{T}_{\text{Dynamic}}$.
  • Static Frame Token. The static frame tokens are extracted from the static frames. Assuming $N_{s}$ static frames are chosen in the video, we uniformly select them from $\bm{F}$. The temporal indices of these frames are:
    =$\begin{gathered}\mathcal{I}=\left\{\left\lfloor\frac{kT}{N_{s}}\right\rfloor+1\mid k=0,1,\dots,N_{s}-1\right\},\end{gathered}$ (1)
    where the first frame is always selected as a static frame. We preserve all tokens from these static frames to constitute $\mathcal{T}_{\text{Static}}$ as:
    $\{\bm{F}^{i,h,w}\mid i\in\mathcal{I},h\in[1,H],w\in[1,W]\},$ (2)
    where $\bm{F}^{i,h,w}\in\mathbb{R}^{D}$ represents the token embedding at the $i$-th frame and spatial location $(h,w)$ in $\bm{F}$. Obviously The total number of static tokens satisfies $|\mathcal{T}_{\text{Static}}|=N_{s}\times H\times W$.
  • Dynamic Delta Token. As illustrated in Figure 3, the static frames divide the entire sampled frame sequence into $N_{s}$ intervals. The frames within these intervals are defined as dynamic frames, and the dynamic delta tokens $\mathcal{T}_{\text{Dynamic}}$ are extracted from them to capture the dynamic information in the video. Since the subsequence length is $L_{b}$, it is evident that the number of dynamic delta tokens satisfies $|\mathcal{T}_{\text{Dynamic}}|=L_{b}-|\mathcal{T}_{\text{Static}}|$. These tokens are uniformly distributed across each interval, meaning the number of dynamic delta tokens selected from the $i$-th interval satisfy $|\mathcal{T}_{\text{Dynamic},i}|=\lfloor|\mathcal{T}_{\text{Dynamic}}|/N_{s}\rfloor$.
    Inspired by video codec, to retain as much dynamic information as possible, we select tokens from each interval that exhibit the largest differences compared to the corresponding tokens in the preceding static frame. We first define the token difference metric based on the cosine similarity between two token embeddings:
    $\mathcal{D}(\bm{f}_{i},\bm{f}_{j})=1-\frac{\bm{f}_{i}\cdot\bm{f}_{j}}{\|\bm{f}_{i}\|\|\bm{f}_{j}\|}.$ (3)
    This metric increases as the two embeddings become more dissimilar. Then, we select $|\mathcal{T}_{\text{Dynamic},i}|$ tokens in the interval $i$ that have the largest differences compared to the corresponding tokens in the preceding static frame. We define $\mathcal{T}_{\text{Dynamic},i}$ as:
    $\begin{gathered}\{\bm{F}_{i}^{j,h,w}\mid\text{Top}_{|\mathcal{T}_{\text{Dynamic},i}|}\mathcal{D}(\bm{F}_{i}^{0,h,w},\bm{F}_{i}^{j,h,w}),\\ j\in[1,T_{i}],h\in[1,H],w\in[1,W]\},\end{gathered}$ (4)
    where $\bm{F}_{i}^{0,h,w}$ denotes the token embedding at position $(h,w)$ in the preceding static frame of current interval $i$, and $\bm{F}_{i}^{j,h,w}$ denotes the token embedding at position $(h,w)$ in the $j$-th dynamic frame of interval $i$. $T_{i}$ refers to the number of dynamic frames in $i$-th interval.
  • Token Subsequence Generation. Then we merge $\mathcal{T}_{\text{Static}}$ and $\mathcal{T}_{\text{Dynamic}}$ according to their original order to obtain the $L_{b}$-long token subsequence $\bm{\hat{T}}_{v}$.
To generate $n$ candidate token subsequences, we vary $N_{s}$ from $1$ to $n$. As $N_{s}$ decreases, the number of dynamic delta tokens increases, thereby preserving more dynamic information under the same token budget. Conversely, the number of static frame tokens increases, preserving more static information. In this way, we can generate token subsequences with varying proportions of static and dynamic information to adapt to the requirements of different queries.
Figure 3: An example of token subsequence generation with 6 total frames and 2 static frames.
4.4 Selection: Quick Evaluation
After obtaining $n$ token subsequences of length $L_{b}$, we perform an evaluation and select the optimal subsequence based on the chosen metric.
Previous studies have identified certain characteristics of visual tokens in attention mechanisms. For instance, Chen et al. (2024a) shows that most vision tokens can be removed at the second layer without significant performance loss, and Wan et al. (2024) observes that vision tokens are generally less attended. Based on these findings, we assume that the attention mechanism at the second layer already provides meaningful clues of token importance. Besides, we hypothesize that inputs with more tokens being attended are more likely to yield better outputs.
To enable quick evaluation, we compute the attention score matrix $\bm{S}$ at the second layer of the VideoLMs, using textual query tokens as the query input, and instruction and vision tokens as the key input:
$\bm{Q}=\bm{W}_{\text{Q}}\bm{H}_{q},$ (5)
$\bm{K}=\bm{W}_{\text{K}}\ \cdot\text{Concat}(\bm{H}_{i},\bm{H}_{v}),$ (6)
$\bm{S}=\text{softmax}\left(\frac{\bm{QK}^{\top}}{\sqrt{d_{k}}}\right),$ (7)
where $\bm{H}_{q}$, $\bm{H}_{i}$, and $\bm{H}_{v}$ denote the hidden features of the textual query, instruction, and visual inputs. And $d_{k}$ is the dimension of key vectors in the attention mechanism. To quantify the attention allocated to visual tokens, we compute the summation of attention scores of the visual tokens. Specifically, to ensure comprehensive consideration of each text query token, we first extract the maximum values along the query dimension from $\bm{S}$, yielding an attention score vector $\bm{s}$ for each visual token. Then we sum the attention scores of the visual tokens:
$\bm{s}=\max_{i}\bm{S}_{ij},$ (8)
$s=\sum_{j=N_{i}}^{N_{i}+N_{v}}\bm{s}_{j},$ (9)
where $\bm{S}_{ij}$ represents the attention score of the $i$-th query token to the $j$-th visual token, $N_{i}$ denotes the number of instruction tokens, and $N_{v}$ is the number of visual tokens. Finally, from the $n$ candidates, we select the input with the highest sum of visual token attention scores as the optimal input:
${\bm{\bar{T}}_{v}}=\underset{m\in\{1,2,\dots,n\}}{\arg\max}s^{(m)},$ (10)
where $s^{(m)}$ denotes the summed attention score for the $m$-th token subsequence.
5 Experiments
5.1 Experiment Settings
Model Settings EgoSchema VideoMME
Method Sample Budget Short Medium Long Overall
Qwen2-VL-7B Original 64 - 66.2 71.1 59.4 50.8 60.4
Retrieval 256 64 63.6 71.0 61.3 52.2 61.5
Similarity 256 64 66.6 71.4 60.6 51.8 61.3
Ours 256 64 67.8 72.4 63.1 53.2 62.9
Original 32 - 64.7 68.9 55.2 48.7 57.6
Retrieval 128 32 61.7 70.0 58.6 51.6 60.0
Similarity 128 32 65.6 70.1 58.7 51.8 60.2
Ours 128 32 66.7 71.4 61.0 51.7 61.4
LLaVA-OneVision-7B Original 64 - 60.1 70.6 55.8 47.8 58.0
Retrieval 256 64 57.7 64.0 53.4 47.0 54.8
Similarity 256 64 59.6 71.0 57.9 50.8 59.9
Ours 256 64 60.3 71.9 58.3 51.4 60.6
Original 32 - 60.4 71.3 57.4 48.0 58.9
Retrieval 128 32 57.9 63.2 53.9 46.0 54.4
Similarity 128 32 60.2 70.8 57.1 49.7 59.2
Ours 128 32 60.5 70.2 58.0 51.6 59.9
Table 2: Long video benchmarks results.
  • Benchmarks. To comprehensively evaluate performance, we select benchmarks for both long and short videos. For long videos, we utilize VideoMME and EgoSchema. And for short videos, we employ MSVD-QA.
    • VideoMME Fu et al. (2024a) comprises 900 videos and 2,700 multiple-choice questions across six domains. Videos range from 11 seconds to 1 hour in duration and are categorized as short, medium, and long subsets.
    • EgoSchema Mangalam et al. (2023) features over 5,000 multiple-choice questions. The videos have an average length of about three minutes and cover a wide range of human activities and behaviors. We use its full set.
    • MSVD-QA Xu et al. (2017) contains 1,970 video clips and about 50,500 question-answer pairs. The average video duration is 10 seconds. We used the test set which includes around 13,000 questions.
    We use multiple-choice accuracy as the metric for VideoMME and EgoSchema. For MSVD-QA, which consists of open-ended questions, we use GPT OpenAI (2024) for evaluation and scoring.
  • Baselines. We validate our method on the currently well recognized Qwen2-VL Wang et al. (2024) and LLaVA-OneVision Li et al. (2024a). Qwen2-VL introduces dynamic resolution support and multimodal rotary position embeddings. LLaVA-OneVision is capable of handling single-image, multi-image, and video tasks. Here we adopt the 7B version for both models.
    As shown in Table 1, previous methods either compress within the KV cache, failing to address the issue of long token sequences input to the large language model, or introduce new modules that require training the model, making it unfair to compare them directly with our training-free approach. Thus we select three baselines: 1) Original: uniform sampling frames within the budget; 2) Retrieval: higher-frequency uniform sampling frames beyond the budget, followed by pruning based on cosine similarity with query embeddings to meet the token limit; 3) Similarity: higher-frequency uniform sampling frames beyond the budget, followed by pruning based on the cosine similarity of adjacent token embeddings to meet the token limit. In practice, the idea of “Retrieval” and “Similarity” are often utilized in compression modules Qian et al. (2025); Song et al. (2024); He et al. (2024a).
    To further demonstrate the advantage of our method, we compare it with 4) LongVU Shen et al. (2024), a training-based approach. For fairness, we reproduce its compression method on both models and evaluate it in a training-free setting.
  • Implementation Details. All experiments are conducted on two 40GB A100 GPUs. For multiple-choice questions, the model generates one token, while for open-ended questions, outputs are limited to 30 tokens. The prompts used for each dataset are detailed in Appendix A.
    Note that video resolution affects the number of frame tokens generated by Qwen2-VL. A fixed token budget may result in inconsistent numbers of frames that can be processed across videos, complicating experiments and comparisons. To address this, we convert the token limit into a frame-based constraint, setting the budget frame number as $T_{b}$. For a video generating $H\times W$ tokens per frame, the token budget becomes $L_{b}=T_{b}\times H\times W$. This adjustment simplifies implementation and ensures fair sample comparisons.
5.2 Main Results
  • Long Video Results. Table 2 presents the results on EgoSchema and VideoMME under two settings: 256-frame sampling with 64-frame budget (256-64) and 128-frame sampling with 32-frame budget (128-32). Our method outperforms baselines across both benchmarks and most subsets.
    Qwen2-VL-7B shows significant improvements under the 256-64 setting, surpassing baselines by up to 4.2% on EgoSchema and 2.5% on VideoMME. In the 128-32 setting, it outperforms baselines by up to 5.0% and 3.8%, and achieves a 5.8% improvement over the “Original” method on VideoMME’s medium subset. LLaVa-OneVision-7B achieves 60.3% and 60.6% accuracy on the two benchmarks under the 256-64 setting, and 60.5% and 59.9% under 128-32, showing weaker advantages. This is likely due to noise introduced by sparse token selection under one-dimensional positional encoding. Qwen2-VL-7B, with its three-dimensional positional encoding, exhibits lower noise, highlighting the benefits of specialized positional embedding design for VideoLMs.
    Overall, these results demonstrate the effectiveness of our method, and reveal some model-specific behaviors and limitations.
  • Short Video Results. Table 3 presents the results on MSVD-QA. Given that average video length is only 10 seconds, we reduce the frame sampling rate to 64 frames and adjust the budget to 16 frames. Our method still achieves good performance on Qwen2-VL-7B, surpassing all baselines. It outperforms the “Retrieval” approach by 2.4% in accuracy and nearly 0.1 in score. However, a contrasting trend is observed on LLaVa-OneVision-7B. While our method exceeds both the “Original” and “Similarity” approach, it is slightly outperformed by the “Retrieval” method, with marginal differences. These results highlight that the baselines exhibit instability across different models and datasets, while our approach delivers robust performance, demonstrating its reliability and effectiveness across diverse experimental settings.
    Model Method MSVD-QA
       Acc Score
    Qwen2-VL Original 66.0 3.59
    Retrieval 64.4 3.52
    Similarity 66.5 3.60
    Ours 66.8 3.61
    LLaVA-OneVision Original 54.3 3.09
    Retrieval 54.8 3.12
    Similarity 54.3 3.10
    Ours 54.7 3.11
    Table 3: Short video benchmark results with 64-frame sampling and 16-frame budget.
    Figure 4: In our qualitative analysis cases, our method allocates static and dynamic tokens at a ratio of 3:1 for Question 1 which is a OCR problem, and a ratio of 5:11 for Question 2 which pertains to action recognition.
    Model Method EgoSchema VideoMME
    Qwen2-VL Original 66.2 60.4
    LongVU 67.2 62.3
    Ours 67.8 62.9
    LLaVA- OneVision Original 60.1 58.0
    LongVU 60.3 59.3
    Ours 60.3 60.6
    Table 4: Comparison with LongVU under a training-free setting with 256-frame sampling and 64-frame budget.
    Method EgoSchema VideoMME
    Orignal 64.7 57.6
    Explore + Random 66.3 60.7
    Explore + Select 66.7 61.4
    Table 5: Ablation about exploration and selection stage.
    Figure 5: Search space size analysis. The x-axis represents the search space size. There are $n$ subsequences in the space, and thier static frame number ranges from $\{1,2,\dots,n\}$. Assuming the budget frame is $N_{b}$, “$1$” refers to $n=N_{b}$, “$1/2$” indicates $n=\lfloor N_{b}/2\rfloor$, “$1/4$” represents $n=\lfloor N_{b}/4\rfloor$.
  • Comparison with training-free LongVU. Many visual token compression methods require training, such as LongVU Shen et al. (2024). For a fair comparison, we implement its compression component in a training-free manner. Following the original paper, we use DINOv2 Oquab et al. (2023) with a 0.83 threshold for frame reduction, downsampling at a $\lfloor 2/3\rfloor$ ratio. In the final step, we find it difficult to remove tokens to meet budget using the threshold. Therefore, we opt for top-K selection for precise token control. As shown in Table 4, our method outperforms LongVU across models and datasets. Moreover, compared to LongVU, our method offers better control over the token count.
5.3 Ablation Study
As shown in Table 5, we conduct a two-stage ablation study on our method. The ablation experiments were performed on Qwen2-VL-7B, sampling 128 frames with a budget of 32 frames. First, we validated the effectiveness of the exploration stage. As indicated by the “Explore” row in the table, generating multiple token subsequences followed by random selection results in improvement compared to the original operation, demonstrating the rationality of our search space design. Then we verify the effectiveness of the selection phase. On both benchmarks, our selection method achieves improvement over the random selection.
5.4 Further Analysis
  • Qualitative Analysis. Figure 4 shows two questions for the same video, which are correctly answered employing our method. Question 1, an OCR problem predominantly reliant on static information, prompts the method to allocate a static-to-dynamic token ratio of 3:1. Conversely, action-ralated Question 2, necessitating the identification of a player scoring a goal, leads to the adoption of a static-to-dynamic token ratio of 5:11.
  • Search Space Size. We study search space size impact using Qwen2-VL-7B, sampling 128 frames with a 32-frame budget. Figure 5 shows performance improves on EgoSchema when search space matches the budget but declines on VideoMME. We attribute this to excessive static frames, causing sparse dynamic delta token selection and deviation from the training distribution, reducing effectiveness. Additionaly, time cost rises with search space size. To balance these, we set the search space to half the budget frame number.
  • Time Overhead. The compression inevitably incurs time overhead, but our focus is on memory efficiency and information retention. Our overhead mainly comes from metric computation. Using shallow attention layers with question-length queries, sequential processing of 16 candidates in a 28-layer model doubles first-token latency but leaves subsequent decoding unaffected, benefiting open-ended questions. Parallel processing can further reduce latency. On LLaVa-OneVision-7B (128-frame sampling with a 32-frame budget), compared to the 2.24s overhead incurred by LongVU, our approach costs only 0.43s without flash-attention.
6 Conclusion
Given that long videos possess tokens far exceeding the capacity that models can process, we advance token compression strategies by unveiling the following crucial fact: different question types exhibit varying dependencies on dynamic and static information. Based on this discovery, we propose a novel token selection strategy for video token compression. Our method splits video frames into static and dynamic categories, and adaptively determines the optimal token allocations among static and dynamic frames guided by each specific query. Experiments demonstrate the effectiveness and generalizability of our method across multiple models and datasets.
Limitations
In this paper, we propose a novel token selection strategy for token compression in video question answering tasks, addressing the varying dependencies of questions on dynamic and static video information. While the effectiveness of our method has been validated across multiple datasets, certain limitations remain. Firstly, due to differences in positional encoding mechanisms across models, some encoding schemes may impact the model’s ability to accurately judge video length and temporally localize events. Nevertheless, we believe our approach holds insight for developing compression modules in pre-trained and fine-tuned video models. Additionally, although our method incurs no additional memory overhead (superior to pruning in the key-value cache), it does introduce time overhead. This overhead mainly stems from metric computation. We utilize the output of a shallow (second-layer) attention mechanism to compute the metric, where only the attention map between query tokens and vision tokens is computed. This overhead only happens during the initial token inference and does not affect subsequent token generation. It is worth noting that such additional time cost is a common challenge for most compression methods.
Appendix A Prompt Details
We utilize the template provided by the model for the instruction prompt part. We only introduce the textual organization format in the questioning part.
A.1 Prompts for Multiple-Choice Questions
We add the sentence "Respond with only the letter (A, B, C, or D) of the correct option." at the beginning of the multiple-choice questions. Here is an example for questions in VideoMME:
Respond with only the letter (A, B, C, or D) of the correct option.
Which elements are depicted in the painting introduced by the video?
A. A little girl and a red balloon.
B. A little boy and a red balloon.
C. A little girl and a blue balloon.
D. An adult and a blue balloon.
And here is an example for EgoSchema:
Respond with only the letter (A, B, C, D or E) of the correct option.
Identify the recurring actions in the video and briefly discuss their significance to the overall narrative.
A. C constantly organizing a plastic box, suggesting her obsession with tidiness
B. C and the boy taking turns throwing objects out of the window, showcasing a game
C. C pouring water and conversing with the boy, highlighting routine and communication
D. The boy trying to get C’s attention by throwing a toy on the blanket repeatedly
E. C teaching the boy how to fold blankets properly and arrange his toys
A.2 Prompts for Open-Ended Questions
We add the sentence "Answer the question according to the video." at the beginning of the multiple-choice questions. Here is an example for questions in MSVD-QA:
Answer the question according to the video.
who did circles on the back tire of his motorcycle in the parking lot?